# Like Father, Like Son: Kinship-Aware Preference Mapping (KARMA) for Automatic Alignment in Large Language Models

### v1

@@ 4.1 RQ1: Dataset Reconstruction using KARMA
To analyze the relationships between models, we computed kinship scores by normalizing evaluation results across six benchmark tasks. Figure 4 presents a Principal Component Analysis (PCA)Maćkiewicz and Ratajczak (1993) visualization of these relationships, illustrating distinct clustering patterns among models. Generally, smaller-scale models tend to cluster in the first quadrant, models with closer kinship relationships in the second quadrant, while Llama-based models and models exceeding 10B parameters are predominantly distributed in the third and fourth quadrants. This distribution underscores the critical role of model training data, training algorithms, and scale in determining kinship relationships among models.
@@
These findings highlight the effectiveness of the KARMA framework in reconstructing preference datasets, providing a more structured and informative approach to model alignment.
4.2 RQ2: Performance of KARMA Binarization
To assess the effectiveness of KARMA binarization, we examined whether model performance can serve as an indicator of data quality. KARMA performs binarization in two sequential steps: first, it defines preference pairs based on kinship between models, and second, it ranks these pairs according to their relative performance. Through this process, we aimed to determine whether response similarity—estimated through model similarity—plays a more significant role in preference modeling than simply prioritizing higher-performing models by consistently positioning them as dominant within each pair.
@@ 4.2 RQ2: Performance of KARMA Binarization
The results of this comparison are presented in Figure 5. Across all benchmark tasks, the kinship-based scoring metric, , demonstrates performance levels nearly equivalent to the existing instance-wise RLAIF method, with a minimal deviation of only 0.008. Notably, for MMLU-Pro and IFeval, KARMA-based binarization even surpasses the RLAIF baseline by approximately 0.01, indicating that structured preference mapping via kinship can yield competitive or superior alignment outcomes. The detailed results, including average scores for each methodology, are summarized in Appendix D.
4.3 RQ3: Model Size and Kinship
The impact of different kinship approaches varies significantly with model size, influencing how models learn from preference data. To investigate this relationship, we conducted experiments using models from the same family but with different parameter counts, specifically comparing small models (3B parameters) and larger models (7-8B parameters). Our analysis reveals a clear pattern in how model size determines the optimal kinship strategy.
Smaller models (3B) exhibit superior performance when trained using the algorithm, which prioritizes learning from response quality within the binarized dataset. In contrast, larger models (7-8B) achieve better results with the algorithm, suggesting that response similarity becomes increasingly important as model size grows.
4.4 RQ4: Model Capability and Kinship Patterns
@@
Models exhibit distinct similarity patterns based on their capabilities, with these patterns varying across different evaluation metrics. Notably, model kinship is influenced by factors such as model scale and training methodology, leading to variations in clustering behavior across different capability dimensions.
@@
5 Conclusion
This paper introduces KARMA, a kinship-based framework for generating high-quality preference data for language model alignment without requiring instance-wise annotation. KARMA achieves performance comparable to RLAIF while offering greater computational efficiency and surpassing baselines on MMLU-Pro and IFeval, demonstrating the effectiveness of kinship-aware preference mapping. Our findings show that model size significantly affects the optimal kinship strategy. Smaller models (3B) benefit from prioritizing response quality, while larger models (7-8B) perform better when leveraging response similarity. Additionally, kinship patterns vary across tasks, with reasoning clustering by model size, instruction-following by model family, and knowledge-based tasks exhibiting a hierarchical structure. These results highlight the need for size-adaptive preference learning strategies and demonstrate that high-quality preference data can be generated without costly manual annotation. Future work will extend KARMA to cross-family model comparisons and further explore the relationship between model kinship and human preference judgments.
6 Related Work and Background
6.1 Alignment Tuning
Alignment tuning has become a central focus in enhancing LLMs to meet user expectations and ethical standardsKumar et al. (2024). Various preference-based learning techniques, particularly reinforcement learning methods such as Proximal Policy Optimization (PPO)Schulman et al. (2017), Direct Preference Optimization (DPO)Rafailov et al. (2024), and Optimized Reinforcement Preference Optimization (ORPO)Hong et al. (2024), have been developed to facilitate this tuning process. These methods rely on preference datasets, which typically contain pairs of responses generated by models based on given instructions, with each pair ranked according to human or model-based evaluations. Prominent alignment datasets like UltrafeedbackCui et al. (2023), which gathers extensive human feedback, and HH-RLHFBai et al. (2022), which uses human-annotated preferences, provide foundational resources for alignment. To minimize the reliance on labor-intensive data curation, recent automated approaches have been introduced to filter or regenerate preference data based on specific criteria, promoting data consistency and scalability. However, such approaches often lack nuanced control over data quality, as they fail to consider the role of the origin model’s capabilities in shaping alignment effectiveness Shi et al. (2024).
6.2 Limitations in Existing Preference Data Approaches
Effective preference data construction requires a clear, rigorous set of criteria to ensure alignment quality across generated response pairs. Common criteria, including Reasoning, Truthfulness, and Instruction-Following, guide the selection of data that aligns with key ethical and functional standardsCui et al. (2023); Bai et al. (2022). High-performing models, capable of producing responses that meet these standards, are often evaluated using benchmarks like ARCClark et al. (2018), MMLUHendrycks et al. (2020), and the Instruction-Following evalZhou et al. (2023) suite, which assess a model’s factual accuracy, reasoning ability, and compliance with instructions. While these benchmarks provide a foundation for selecting responses that promote alignment goals, existing methods typically overlook the impact of variation in model capabilities on alignment consistency.
The KARMA Framework addresses this limitation by introducing a kinship-aware approach that emphasizes model compatibility in preference data selection. By curating preference pairs based on model kinship—aligning models with similar core competencies in knowledge, reasoning, and instruction-following—the KARMA Framework enhances alignment coherence and ensures stable, high-quality data. Further, its curriculum-based selection process sequences preference data from simpler to more complex distinctions, optimizing the alignment process for improved consistency and scalability.
Limitation
In this paper, we demonstrated that KARMA can achieve performance comparable to existing RLAIF without requiring annotations for SFT and DPO. However, due to resource and time constraints, we were unable to validate our approach across a broader range of alignment tuning techniques. Further evaluation is needed for methods.
Additionally, our evaluation primarily focused on fundamental abilities such as Reasoning, Knowledge, and Instruction-Following. However, we did not assess KARMA’s performance on other important values, including factuality and ethical standard. Future work should incorporate evaluations reflecting these aspects.
Appendix A Implementation Details for Model Training
Appendix B Prompt template for Evaluation Benchmarks
To determine kinship relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results. The benchmark performance of all 17 models considered in this study is presented in Table 8. For the start and end tokens, the template recommended in the paper was used.
Appendix C Evaluation Origin Model Capabilities for Kinship
To determine kinship relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results. The benchmark performance of all 17 models considered in this study is presented in Table 3. The actual calculation example for and can be found in Figure 9.
Appendix D Total Performance
@@
| Benchmark | Model | Baseline | ||||

### v2

@@ 5.1 RQ1: Dataset Reconstruction using ZEBRA
To analyze the relationships between models, we computed SUP and SIM by normalizing evaluation results across six benchmark tasks. Figure 6 presents a Principal Component Analysis (PCA)Maćkiewicz and Ratajczak (1993) visualization of these relationships, illustrating distinct clustering patterns among models. Generally, smaller-scale models tend to cluster in the first quadrant, models with closer Model Behavior relationships in the second quadrant, while Llama-based models and models exceeding 10B parameters are predominantly distributed in the third and fourth quadrants. This distribution underscores the critical role of model training data, training algorithms, and scale in determining Model Behavior relationships among models.
We further verify that anchor pairs selected by SIM indeed yield semantically closer responses. Using a TF–IDFRamos (2003), we obtain an average response-pair similarity of 0.4623 for the most similar model pair () versus 0.4129 for the least similar model pair (). This 12% relative gap confirms that high-MB-SIM anchoring indeed surfaces finer-grained yet coherent preference signals.
@@
These findings highlight the effectiveness of the ZEBRA framework in reconstructing preference datasets, providing a more structured and informative approach to model alignment.
5.2 RQ2: Performance of ZEBRA Binarization
To assess the effectiveness of ZEBRA binarization, we examined whether model performance can serve as an indicator of data quality. The results of this comparison are presented in table 1. Across all benchmark tasks, the Model Behavior-based scoring metric, SUP, demonstrates performance levels nearly equivalent to the instance-wise RLAIF method, with a minimal deviation of only 0.008.
@@ 5.2 RQ2: Performance of ZEBRA Binarization
Notably, for MMLU-Pro and IFeval, ZEBRA-based binarization even surpasses the RLAIF baseline by approximately 0.01, indicating that structured preference mapping via Model Behavior can yield competitive or superior alignment outcomes. Table 1 shows the results of the average scores for each methodology. The detailed results, including average scores for each methodology, are summarized in Appendix E.
Statistical significance.
Table 2 reports paired -tests and permutation tests (10,000 shuffles) that compare each anchoring strategy with the baseline instance-wise RLAIF data across the same 1 024 evaluation prompts. None of the strategies achieves statistical significance at , although the hybrid SUP+SIM shows the largest absolute gain (, , ). These numbers indicate that ZEBRA’s zero-annotation binarization maintains baseline-level alignment quality (within 3 pp) without incurring per-instance scoring cost, even if a decisive improvement is not yet observed.
5.3 RQ3: Cost & Efficiency Analysis
Table 3 contrasts the labeling cost of ZEBRA with canonical RLHF and RLAIF pipelines. The standing out points is below: Absolute cost gap. LLM-annotated RLAIF corpora lower this to $0.6–$4 K by outsourcing each pairwise judgment to GPT-4, yet they still purchase every label. ZEBRA, by reusing benchmark leaderboards, pays no marginal cost ($0) for preference construction. Relative efficiency. Normalised per comparison, GPT-4 labels ($0.063) are cheaper than human labels ($0.67), but ZEBRA is orders of magnitude cheaper than both because it dispenses with pairwise annotation altogether.
Cost matters only if quality survives. Despite a zero-dollar label budget, ZEBRA matches or surpasses RLAIF baseline (Table 1); the mean performance difference across six benchmarks is .
Consequently, ZEBRA offers a cost-minimal, scalable, and annotation-free route to high-quality preference data. Because the price of human labor or GPT-4 tokens scales linearly with data volume, traditional pipelines become progressively more expensive as models, tasks, and safety domains proliferate. ZEBRA decouples alignment from annotation cost: adding a new model needs no further labels, and incorporating an extra benchmark simply augments the behavior matrix.
6 Conclusion
We presented ZEBRA, a zero-annotation framework that replaces per-instance preference labeling with model-behavior knowledge distilled from public benchmark tables. By pairing responses according to either behavioral superiority or similarity—or a hybrid of the two—ZEBRA constructs 64k high-quality preference pairs at zero marginal cost, completely eliminating the dominant expense in RLHF or RLAIF pipelines.
Experiments across 6 standardized benchmarks show that models fine-tuned with SUP achieve performance on par with, and sometimes better than, an RLAIF baseline. SIM and the hybrid SUP+SIM remain within a small, statistically insignificant margin of the baseline. These results confirm that benchmark-derived behavioral knowledge can serve as an effective proxy for preference supervision. Crucially, ZEBRA reduces marginal labeling cost to zero, reducing tens of thousands of dollars otherwise required for large-scale preference annotation.
Future work will explore richer behavioral axes such as safety, toxicity, and multilingual ability; investigate adaptive schedules that incorporate additional refined pairing strategies during training; and evaluate how well ZEBRA preferences correlate with explicit human judgments.
7 Related Work and Background
7.1 Alignment Tuning
Alignment tuning has become a central focus in enhancing LLMs to meet user expectations and ethical standardsKumar et al. (2024). Various preference-based learning techniques, particularly reinforcement learning methodsSchulman et al. (2017); Rafailov et al. (2024); Hong et al. (2024), have been developed to facilitate this tuning process.
These methods rely on preference datasets, which typically contain pairs of responses generated by models based on given instructions, with each pair ranked according to human or model-based evaluations. Prominent alignment datasets like UltrafeedbackCui et al. (2023), which gathers extensive human feedback, and HH-RLHFBai et al. (2022), which uses human-annotated preferences, provide foundational resources for alignment. To minimize the reliance on labor-intensive data curation, recent automated approaches have been introduced to filter or regenerate preference data based on specific criteria, promoting data consistency and scalability. However, such approaches often lack nuanced control over data quality, as they fail to consider the role of the origin model’s capabilities in shaping alignment effectiveness Shi et al. (2024).
7.2 Limitations in Existing Preference Data Approaches
Effective preference data construction requires a clear, rigorous set of criteria to ensure alignment quality across generated response pairs. Common criteria, including Reasoning, Truthfulness, and Instruction-Following, guide the selection of data that aligns with key ethical and functional standardsCui et al. (2023); Bai et al. (2022). High-performing models, capable of producing responses that meet these standards, are often evaluated using benchmarks like ARCClark et al. (2018), MMLUHendrycks et al. (2020), and the Instruction-Following evalZhou et al. (2023) suite, which assess a model’s factual accuracy, reasoning ability, and compliance with instructions.
The ZEBRA Framework addresses this limitation by introducing a model behavior-level approach that emphasizes model compatibility in preference data selection. By curating preference pairs based on model behavior knowledge with similar core competencies in knowledge, reasoning, and instruction-following—the ZEBRA Framework enhances alignment coherence and ensures stable, high-quality data.
Limitation
In this paper, we demonstrated that ZEBRA can achieve performance comparable to existing RLAIF without requiring annotations for SFT and DPO. However, due to resource and time constraints, we were unable to validate our approach across a broader range of alignment tuning techniques. Further evaluation is needed for methods.
Additionally, our evaluation primarily focused on fundamental abilities such as Reasoning, Knowledge, and Instruction-Following. However, we did not assess ZEBRA’s performance on other important values, including factuality and ethical standard. Future work should incorporate evaluations reflecting these aspects.
Appendix A Implementation Details for Model Training
Appendix B Prompt template for Evaluation Benchmarks
To determine behavior knowledge between models, this paper evaluates benchmark performance and compares the similarity of these numerical results. The benchmark performance of all 17 models considered in this study is presented in Table 7. For the start and end tokens, the template recommended in the paper was used.
Appendix C Evaluation Model Benchmark Performance
To determine behavior knowledge relationships between models, this paper evaluates benchmark performance and compares the similarity of these numerical results. The benchmark performance of all 17 models considered in this study is presented in Table 4. The actual calculation example for and can be found in Figure 10.
Appendix D Benchmark Performance Representation
@@
What benchmark represent the model capability: knowledge, instruction-following, and reasoning? Models exhibit distinct similarity patterns based on their capabilities, with these patterns varying across different evaluation metrics. Notably, model behavior knowledge is influenced by factors such as model scale and training methodology, leading to variations in clustering behavior across different capability dimensions.
@@
Figure 8 shows the frequency of each model being selected as the positive or negative under each strategy. Overall, larger and more familiar models tend to be chosen more frequently.
Appendix E Total Performance
@@
Appendix F Model Size and ZEBRA Strategies
The impact of different Strategies varies significantly with model size, influencing how models learn from preference data. To investigate this relationship, we conducted experiments using models from the same family but with different parameter counts, specifically comparing small models (3B parameters) and larger models (7-8B parameters). Our analysis reveals a clear pattern in how model size determines the optimal strategy.
Smaller models (3B) exhibit superior performance when trained using the SUP algorithm, which prioritizes learning from response quality within the binarized dataset. In contrast, larger models (7-8B) achieve better results with the SIM algorithm, suggesting that response similarity becomes increasingly important as model size grows.
| Benchmark | Model | Baseline | SUP | SIM | SUP+SIM | |